# Resonance of an object floating within a surface wavefield

### v2

@@ IV CONCLUSION
We have investigated the interaction of a floating cylinder with surface waves in laboratory-scale experiments. In the first part, the bodies’ natural oscillation frequencies were experimentally characterized and theoretically rationalized without external wave fields. We show that the resonance frequency of the heaving motion corresponds to the maximum wave emission. The resonance frequency $f_{0}$ depends on the cylinder radius $R$ and its thickness $h$. We present a theoretical model that rationalizes the observed resonance frequencies for all the body dimensions characterized experimentally. We show that, in the high-frequency limit, the added mass of a floating body corresponds to half the same body’s added mass when fully submerged in the fluid bulk. Although this hypothesis is not fulfilled for heaving motions, the comparison with the experimental resonant frequency is quantitative. In the second part, we studied the effects of the floater’s vertical motion on the re-emitted wave field. We measured the wave field re-emitted by the floating cylinder and showed that, in the transverse direction, the re-emitted wave exhibits a minimum emission for an excitation frequency close to the body resonance frequency $f_{0}$. By comparing the wave fields re-emitted by a fixed and a free-floating cylinder with the same dimension, we proved that the body oscillations are indeed responsible for the observed minimal transverse emission.
@@ Acknowledgements.
This work has benefited from the financial support of Mairie de Paris through Emergence(s) grant 2021-DAE-100 245973, and the Agence Nationale de la Recherche through grant MSIM ANR-23-CE01-0020-02. W.R. is supported by the Italian “Ministero dell’Università e della Ricerca” (MUR) through the program PON “Ricerca e Innovazione” 2014–2020.
@@
I. The Disk Limit: An Uniform Approximation
To study the case of a thin floating disk, we can take the limit $\epsilon\rightarrow 0$. In the expression of the resonance frequency of the main text, we see that the product $\epsilon\alpha(\epsilon)$ tends to $2/\pi$. Thus, for small values of $\epsilon$, the frequency reduces to
@@ I. The Disk Limit: An Uniform Approximation
$10.96<\Omega(\epsilon=0)/\Omega(\epsilon)<$ 1

### v3

@@ IV CONCLUSION
We have investigated the interaction of a floating cylinder with surface waves in laboratory-scale experiments. In the first part, the bodies’ natural oscillation frequencies were experimentally characterized thanks to a demodulation method which enabled the reconstruction of the surface deformation with a high resolution in both space and time. We show that the resonance frequency of the heaving motion corresponds to a maximum wave emission. The resonance frequency $f_{0}$ depends on the cylinder radius $R$ and its thickness $h$. We present a theoretical model that rationalizes the observed resonance frequencies for all the body dimensions characterized experimentally. We show that, in the high-frequency limit, the added mass of a floating body corresponds to half the same body’s added mass when fully submerged in the fluid bulk. Although the high-frequency hypothesis is not strictly fulfilled for the heaving motions in our experiments, the comparison with the experimental resonant frequency is quantitative. In the second part, we studied the effects of the floater’s vertical motion on the re-emitted wave field. We measured the wave field re-emitted by the floating cylinder and showed that, in the transverse direction, the re-emitted wave exhibits a minimum emission for an excitation frequency close to the body resonance frequency $f_{0}$. By comparing the wave fields re-emitted by a fixed and a free-floating cylinder with the same dimension, we proved that the body oscillations are indeed responsible for the observed minimal transverse emission.
@@ Acknowledgements.
This work has benefited from the financial support of Mairie de Paris through Emergence(s) grant 2021-DAE-100 245973, and the Agence Nationale de la Recherche through grant MSIM ANR-23-CE01-0020-02. W.R. is supported by the Italian Ministero dell’Università e della Ricerca (MUR) through the program PON Ricerca e Innovazione 2014-2020.
@@
I. The Disk Limit: A Uniform Approximation
To study the case of a thin floating disk, we can take the limit $\epsilon\rightarrow 0$. In the expression of the resonant frequency of the main text, we see that the product $\epsilon\alpha(\epsilon)$ tends to $2/\pi$. Thus, for small values of $\epsilon$, the frequency reduces to
@@ I. The Disk Limit: A Uniform Approximation
$0.96<\Omega(\epsilon=0)/\Omega(\epsilon)<1$